A Disc Of mass m, radius r is given an angular velocity w0(omega not). it is then placed on a horizontal rough surface with coeffecient of friction u(myu).


Find the time in which it stops.

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

areal mass density=m/(pi*r^2)=rho

mass of circular loop at radius a is rho*2*(pi)rdr=dm

Torque on the loop=(myu)*g*dm*a

Total torque=(myu)g*∫a dm (from a=0 to r)

angular deceleration=(torque/moment of inertia)

w0=(angular deceleration)*(time in which it stops)